silvernight269:

Write the equation of the polynomial (of lowest possible degree) which needs the given conditions. Write the equation in factored for. Has a local minimum at the point (-2,0) and a triple zero at x=3

Nnesha:

given: 3 zeros ( x-intercept when y=0) use the equation \[\rm f(x)=a(x-z_1)(x-z_2)(x-z_3)\] to find polynomial where (x-z) are the factors. x=3 subtract 3 from both side to write it as a factor we will get `x-3=0` since there are triple zeros at x=3 the equation will be \[\rm f(x)=a(x-3)(x-3)(x-3)\] f(x) is same as y plugin the given points to solve for a